OpenStudy (anonymous):

completing the square.- x^2+3x-18=0 so..x^2+3x_=18+_..... _ = ?

OpenStudy (anonymous):

why -6,3 ?!

jimthompson5910 (jim_thompson5910):

Is the expression \[\large - x^2+3x-18=0\] or is it \[\large x^2+3x-18=0\] ???

OpenStudy (anonymous):

2nd one

jimthompson5910 (jim_thompson5910):

\[\large x^2+3x-18=0\] \[\large x^2+3x=18\] \[\large x^2+3x+\frac{9}{4}=18+\frac{9}{4}\] \[\large (x+\frac{3}{2})^2=\frac{81}{4}\] \[\large x+\frac{3}{2}=\pm\sqrt{\frac{81}{4}}\] \[\large x+\frac{3}{2}=\pm\frac{9}{2}\] \[\large x=-\frac{3}{2}\pm\frac{9}{2}\] \[\large x=-\frac{3}{2}+\frac{9}{2} \ \text{or} \ \ x=-\frac{3}{2}-\frac{9}{2}\] \[\large x=\frac{6}{2} \ \text{or} \ \ x=-\frac{12}{2}\] \[\large x=3\ \text{or} \ \ x=-6\]

OpenStudy (anonymous):

Thanks a lot but...where in hell did you got the 9/4 from?

jimthompson5910 (jim_thompson5910):

I took half of the x coefficient 3 to get \[\frac32\] then I squared that fraction to get \[\frac94\]. I then added this to both sides

jimthompson5910 (jim_thompson5910):

This is part of what it takes to complete the square

OpenStudy (anonymous):

thanks

jimthompson5910 (jim_thompson5910):

np